OpenStudy (anonymous):

In 1984, the Fizzy Cola Company sold 23 million gallons of soda. by 2003, the company was selling 127 million gallons of soda. What is the average rate of change in the number of gallons of soda per year? HELP!!

OpenStudy (anonymous):

Increase in gallons divided by number of years: Increase in gallons= 127M-23M=104M Number of years=2003-1984=19

OpenStudy (anonymous):

so 104 divided by 19?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

okay thanks!! :)

OpenStudy (anonymous):

You can get the equation for any year in millions of gallons: \[G=\frac{ 104 }{ 19 }(Y-1984)+23\]

OpenStudy (anonymous):

awesome thanks so much!!

OpenStudy (anonymous):

welcome!